# Coupler enabled tunable dipole-dipole coupling between optically levitated nanoparticles

### v1

@@
Multiple optically levitated particles in vacuum can exhibit electrostatic interactions, optical binding, or non-reciprocal light-induced dipole-dipole interactions, making them promising platforms for exploring mesoscopic entanglement and complex interactions. However, in optical trap arrays, individually controlling the position and polarization of each trap is challenging, limiting the precise tuning of interactions between adjacent particles. This constraint hinders the study of complex interaction systems. In this work, we introduce a third nanoparticle as a coupler to two initially non-interacting nanoparticles, achieving tunable dipole-dipole coupling mediated by the third one. We investigated the effect of the particles’ phases and positions on the interaction strength and demonstrated its broad tunability. Our method allows for precise control of interactions between any pair of adjacent particles in multi-particle systems, facilitating the further use of levitated nanoparticle arrays in entanglement research and sensing.
@@ 1 Introduction
Similar to optical tweezer arrays of levitated microscopic atoms or molecules, arrays of levitated macroscopic nanoparticles [25] with tunable dipole-dipole coupling are recognized as essential for achieving macroscopic entanglement [14, 16] and exotic topological quantum states [26, 27]. Interestingly, the coupling between optically levitated nanoparticles can be Hermitian or non-Hermitian due to scattering interactions, introducing a wide range of phenomena, including limit cycles [28, 29] and other exotic behaviors [30, 31, 32, 33, 34, 35].
@@
Here we demonstrate that an ancillary silica nanoparticle can mediate the coupling between the dipole-dipole interactions between other two nanoparticles, without requiring any inherent interaction between the two particles, as shown in Fig. 1(a). We investigate how the dipole-dipole interactions mediated by the third particle vary with the phase difference, interparticle distances, and relative positions between the three particles, demonstrating the tunability of the interaction strength. Our work provides a novel approach for precisely controlling interactions between adjacent particles in multi-body systems, bypassing the problem of untunable position and polarization of each optical trap in optical trap arrays. This will facilitate the construction of larger-scale, fully programmable mechanical oscillator arrays and further applications in macroscopic quantum mechanics [13, 14, 15] and stochastic thermodynamics [38].
Figure 1: (a) Schematic of light-induced dipole-dipole interactions between two optically levitated nanoparticles mediated by a third one. The blue particles represent a one-dimensional optical trap array composed of multiple particles, while the orange particle represents a particle acting as a coupler to mediate interactions between any pair of adjacent particles in the optical trap array. The red arrows represent light-induced dipole-dipole interactions between the particles. (b) In the focal plane, particle 1, particle 2, and particle 3 are trapped in trap 1, trap 2, and trap 3, respectively. Trap 1 and trap 3 are distributed along the $x$ axis, with their polarization aligned along the $x$ axis. Trap 2 is located on the perpendicular bisector of the line segment connecting trap 1 and trap 3, and its polarization is parallel to the $y$ axis. The angle between the line connecting trap 1 and trap 2 and the $x$ axis is $\theta$, and the same applies to the line connecting trap 2 and trap 3. The spacing between trap 1 and trap 3 is $d$. The figure also illustrates the characteristic spatial profiles of the dipole radiation in the far field. (c) Camera image of three nanoparticles trapped in three optical traps under 532 nm laser illumination ($\theta\approx 45^{\circ}$, $d\approx 4.24\ \mu\rm{m}$).
@@ 2 Theoretical model
Here, $G$ is a constant that is positively related to the trapping power and the particle polarizability. Considering the phase difference $\pi-\Delta\varphi$ between particle 2 and particle 3, the coupling constant from particle 3 to particle 2 is $-k_{1}+k_{2}$, and from particle 2 to particle 3 is $-k_{1}-k_{2}$. We can now derive the linear dynamics equations for the system composed of the three optically levitated particles:
@@
Here, $z_{i}$ represents the coordinate of the particle $i$ along the optical axis, and the three particles share the same mass $m$. Furthermore, we assume ${\Delta\varphi}=0$, thereby setting $k_{2}=0$ and turning the non-reciprocal interactions off. Then we can obtain three eigenfrequencies $\omega_{1,2,3}$ and the corresponding eigenvectors of the coupled harmonic oscillators by solving Eqs. (3). It is worth mentioning that in our previous experiment, we measured the coupling coefficient in the presence of particle 1 and particle 2. When the common intrinsic mechanical frequency of two particles was $\Omega\approx 47\cdot 2\pi\ \rm{kHz}$, the angle was $\theta=45^{\circ}$, the phase difference was $\Delta\phi=0$, and the distance was $d=3.11\ \mu\rm{m}$, we measured the value of the conservative coupling coefficient ${k_{1}}/m\Omega\approx 6.0\cdot 2\pi\ \rm{kHz}$. From this value, we can further estimate the conservative coupling coefficient in other cases, which will be used for subsequent simulations.
Figure 2: (a) The eigenfrequencies $\omega_{1,2,3}$ of the coupled oscillators vary with the interparticle distance $d$ at $\theta=45^{\circ}$, ${\Omega}\approx 47\cdot 2\pi\ \rm{kHz}$, $\eta\approx-0.2$, and ${\Delta\varphi}=0$. The curve exhibits a periodicity of $\sqrt{2}\lambda/2$. Addionally, the motion states of particle 1, particle 2, and particle 3 along the $z$ axis in the three normal modes of the system are illustrated in the figure. (b) The eigenfrequencies $\omega_{1,2,3}$ of the coupled oscillators vary with the angle $\theta$ at $d=3.01\ \mu\rm{m}$, ${\Omega}\approx 47\cdot 2\pi\ \rm{kHz}$, $\eta\approx-0.2$, and ${\Delta\varphi}=0$. This figure also shows the distribution diagrams of three optical traps in the $x-y$ plane at angles $\theta=20^{\circ}$, $\theta=45^{\circ}$, and $\theta=70^{\circ}$.
@@ 3 Experimental setup
Figure 3: Schematic of the experimental setup. A 1064 nm trapping beam passes through a beam expander (BE, BE03-1064, Thorlabs) and is incident on a spatial light modulator (SLM), using phase modulation to generate three optical traps. A high-voltage DC power supply ($\rm{HV}\sim 800\rm{V}$) is connected to a pair of bare wires inside the vacuum chamber to change the net charge of the particals through corona discharge. Furthermore, an electric driving signal (ED) is fed into a pair of steel electrodes (SE) to monitor the number of charges by driving the particles. The particles’ signals are synchronously recorded by quadrant photodetectors QPD1 and QPD2. PBS, polarizing beam splitter; HWP, half-wave plate; L, lens; MO, microscope objective; AsL, aspheric lens; M, mirror.
@@ 4 Phase-dependence of interactions
Figure 4: (a) The normal mode splitting, defined as $|\omega_{2}-\omega_{1}|$, varies with the phase $\Delta\varphi$ when $d=3.01\ \mu\rm{m}$, $\theta=45^{\circ}$, and $\eta\approx-0.24$. The purple dots represent the experimentally measured normal mode splitting. The red curve represents the simulation results incorporating the intrinsic mechanical frequencies $\Omega_{1}$, $\Omega_{2}$, and $\Omega_{3}$ under experimental conditions. When the normal mode splitting is smaller than the mechanical linewidth $\gamma$ (indicated by the dashed gray line), the avoided crossing is indiscernible. (b) The PSD of the $z$-mode signal is shown for cases where conservative interactions dominate (blue data, $\Delta\varphi=\pi$) and where non-conservative interactions dominate (red data, $\Delta\varphi=\pi/2$) when $d=2.80\ \mu\rm{m}$, $\theta=45^{\circ}$, and $\eta\approx-0.05$.
@@
In order to further illustrate the effect of non-conservative interactions on the normal modes of the three-particle system, we recorded the PSD of particle 1 and particle 3 at $d=2.80\ \mu\rm{m}$, $\theta=45^{\circ}$, and $\eta\approx-0.05$ (where both conservative and non-conservative interactions are present), as shown in Fig. 4(b). When the phase difference reaches $\Delta\varphi=\pi$, the conservative interactions dominate, resulting in non-degenerate normal modes with typical mode splitting. However, when the phase difference reaches $\Delta\varphi=\pi/2$, the non-conservative interactions dominate, causing two of the normal modes to become degenerate. It results in resonance and a significant amplification in the amplitude of particle motion. In our experiment, the presence of non-conservative interactions makes it difficult to observe normal mode splitting and may even lead to particle escape.
@@ 5 Position-dependence of interactions
Figure 5: (a) The normal mode splitting $|\omega_{2}-\omega_{1}|$ as a function of interparticle distance $d$ is shown for $\theta=45^{\circ}$, $\eta\approx-0.24$, and $\Delta\varphi=0$. The blue dots represent the experimentally measured normal mode splitting, while the red curve depicts the simulation results incorporating the intrinsic mechanical frequencies $\Omega_{1}$, $\Omega_{2}$, and $\Omega_{3}$ for varying distances. The curve exhibits a periodicity of $\sim\sqrt{2}\lambda/2$. The avoided crossing is unresolved when the splitting is smaller than the mechanical linewidth $\gamma$ (indicated by the dashed gray line). (b-e) The PSD of the $z$-mode signals is presented for interparticle distances $d=3.01\ \mu\rm{m}$, $3.43\ \mu\rm{m}$, $3.81\ \mu\rm{m}$, and $4.24\ \mu\rm{m}$ with $\theta=45^{\circ}$, $\eta\approx-0.24$, and $\Delta\varphi=0$. At these distances, the normal modes $\omega_{1}$ and $\omega_{2}$ exhibit variations between maximum splitting and no splitting." (f) The normal mode splitting $|\omega_{2}-\omega_{1}|$ as a function of the angle $\theta$ is shown for $d=3.01\ \mu\text{m}$, $\eta\approx-0.05$, and $\Delta\varphi=0$. The magenta dots represent the experimentally measured normal mode splitting, while the red curve illustrates the simulation results. As the angle $\theta$ varies from $35.4^{\circ}$ to $60.9^{\circ}$, $|\omega_{2}-\omega_{1}|$ oscillates, with both the period and amplitude of the oscillation gradually decreasing. The gray dashed line represents the mechanical linewidth $\gamma$.

### v2

@@
Multiple optically levitated particles in vacuum can exhibit electrostatic interactions, optical binding, and non-reciprocal light-induced dipole-dipole interactions, making them promising platforms for exploring mesoscopic entanglement and complex interactions. However, in optical trap arrays, individually controlling the position and polarization of each trap is challenging, limiting the precise tuning of interactions between adjacent particles. This constraint hinders the study of complex interaction systems. In this work, we introduce a third nanoparticle as a coupler to two initially non-interacting nanoparticles, achieving tunable dipole-dipole coupling mediated by the third one. We investigated the effect of the particles’ phases and positions on the interaction strength and demonstrated its broad tunability. Our method allows for precise control of interactions between any pair of adjacent particles in multi-particle systems, facilitating the further use of levitated nanoparticle arrays in macroscopic quantum mechanics and sensing.
@@ 1 Introduction
Similar to optical tweezer arrays of levitated microscopic atoms or molecules, arrays of levitated macroscopic nanoparticles [25] with tunable dipole-dipole coupling are recognized as essential for achieving macroscopic entanglement in an optical cavity [26] and exotic topological quantum states [27, 28]. Interestingly, the coupling between optically levitated nanoparticles can be Hermitian or non-Hermitian due to scattering interactions, introducing a wide range of phenomena, including limit cycles [29, 30] and other exotic behaviors [31, 32, 33, 34, 35, 36].
@@
Here we demonstrate that an ancillary silica nanoparticle can mediate the dipole-dipole coupling between other two nanoparticles, without requiring any inherent interaction between the two particles, as shown in Fig. 1(a). We investigate how the dipole-dipole interactions mediated by the third particle vary with the phase difference, interparticle distances, and relative positions between the three particles, demonstrating the tunability of the interaction strength. Our work provides a novel approach for precisely controlling interactions between adjacent particles in multi-body systems, bypassing the problem of untunable position and polarization of each optical trap in optical trap arrays. This will facilitate the construction of larger-scale, fully programmable mechanical oscillator arrays and further applications in macroscopic quantum mechanics [13, 14, 15] and stochastic thermodynamics [39].
Figure 1: (a) Schematic of light-induced dipole-dipole interactions between two optically levitated nanoparticles mediated by a third one. The blue particles represent a one-dimensional optical trap array composed of multiple particles, while the orange particle represents a particle acting as a coupler to mediate interactions between any pair of adjacent particles in the optical trap array. The red arrows represent light-induced dipole-dipole interactions between the particles. (b) In the focal plane, particle 1, particle 2, and particle 3 are trapped in trap 1, trap 2, and trap 3, respectively. Trap 1 and trap 3 are distributed along the $x$ axis, with their polarization aligned along the $x$ axis. Trap 2 is located on the perpendicular bisector of the line segment connecting trap 1 and trap 3, and its polarization is parallel to the $y$ axis. The angle between the line connecting trap 1 and trap 2 and the $x$ axis is $\theta$, and the same applies to the line connecting trap 2 and trap 3. The interparticle distance between trap 1 and trap 3 is $d$. The figure also illustrates the characteristic spatial profiles of the dipole radiation in the far field. (c) Camera image of three nanoparticles trapped in three optical traps under 532 nm laser illumination ($\theta\approx 45^{\circ}$, $d\approx 4.24\ \mu\rm{m}$).
@@ 2 Theoretical model
Here, $G$ is a constant that is positively related to the trapping power and the particle polarizability, and $k$ is the wave number of the trapping light. Considering the phase difference $\pi-\Delta\varphi$ between particle 2 and particle 3, the coupling constant from particle 3 to particle 2 is $-k_{1}+k_{2}$, and from particle 2 to particle 3 is $-k_{1}-k_{2}$. We can now derive the linear dynamics equations for the system composed of the three optically levitated particles:
@@
Here, $z_{i}$ denotes the position of particle $i$ along the optical axis, while $\gamma$ represents the mechanical linewidth, which also corresponds to the damping rate. The three particles share the same mass $m$. Furthermore, we assume ${\Delta\varphi}=0$, thereby setting $k_{2}=0$ and turning the non-reciprocal interactions off. Then we can obtain three eigenfrequencies $\omega_{1,2,3}$ and the corresponding eigenvectors of the coupled harmonic oscillators by solving Eqs. (3). It is worth mentioning that in our previous experiment, we measured the coupling coefficient in the presence of particle 1 and particle 2. When the common intrinsic mechanical frequency of two particles was $\Omega\approx 47\cdot 2\pi\ \rm{kHz}$, the angle was $\theta=45^{\circ}$, the phase difference was $\Delta\phi=0$, and the distance was $d=3.11\ \mu\rm{m}$, we measured the value of the conservative coupling coefficient ${k_{1}}/m\Omega\approx 6.0\cdot 2\pi\ \rm{kHz}$. From this value, we can further estimate the conservative coupling coefficient in other cases, which will be used for subsequent simulations.
Figure 2: (a) The eigenfrequencies $\omega_{1,2,3}$ of the coupled oscillators vary with the interparticle distance $d$ at $\theta=45^{\circ}$, ${\Omega}\approx 47\cdot 2\pi\ \rm{kHz}$, $\eta\approx-0.2$, and ${\Delta\varphi}=0$. The curve exhibits a periodicity of $\sqrt{2}\lambda/2$. Additionally, the motion states of particle 1, particle 2, and particle 3 along the $z$ axis in the three normal modes of the system are illustrated in the figure. (b) The eigenfrequencies $\omega_{1,2,3}$ of the coupled oscillators vary with the angle $\theta$ at $d=3.01\ \mu\rm{m}$, ${\Omega}\approx 47\cdot 2\pi\ \rm{kHz}$, $\eta\approx-0.2$, and ${\Delta\varphi}=0$. This figure also shows the distribution diagrams of three optical traps in the $x-y$ plane at angles $\theta=20^{\circ}$, $\theta=45^{\circ}$, and $\theta=70^{\circ}$.
@@ 3 Experimental setup
Figure 3: Schematic of the experimental setup. A 1064 nm trapping beam passes through a beam expander (BE, BE03-1064, Thorlabs) and is incident on a spatial light modulator (SLM), using phase modulation to generate three optical traps. A high-voltage DC power supply ($\rm{HV}\sim 800\rm{V}$) is connected to a pair of bare wires inside the vacuum chamber to change the net charge of the particles through corona discharge. Furthermore, an electric driving signal (ED) is fed into a pair of steel electrodes (SE) to monitor the number of charges by driving the particles. The particles’ signals are synchronously recorded by quadrant photodetectors QPD1 and QPD2. PBS, polarizing beam splitter; HWP, half-wave plate; L, lens; MO, microscope objective; AsL, aspheric lens; M, mirror.
@@ 4 Phase-dependence of interactions
Figure 4: (a) The normal mode splitting, defined as $|\omega_{2}-\omega_{1}|$, varies with the phase difference $\Delta\varphi$ when $d=3.01\ \mu\rm{m}$, $\theta=45^{\circ}$, and $\eta\approx-0.24$. The purple dots represent the experimentally measured normal mode splitting. The red curve represents the simulation results incorporating the intrinsic mechanical frequencies $\Omega_{1}$, $\Omega_{2}$, and $\Omega_{3}$ under experimental conditions. When the normal mode splitting is smaller than the mechanical linewidth $\gamma$ (indicated by the dashed gray line), the avoided crossing is indiscernible. (b) The PSD of the $z$-mode signal is shown for cases where conservative interactions dominate (blue data, $\Delta\varphi=\pi$) and where non-conservative interactions dominate (red data, $\Delta\varphi=1.39\pi$) when $d=2.80\ \mu\rm{m}$, $\theta=45^{\circ}$, and $\eta\approx-0.05$.
@@
In order to further illustrate the effect of non-conservative interactions on the normal modes of the three-particle system, we recorded the PSD of particle 1 and particle 3 at $d=2.80\ \mu\rm{m}$, $\theta=45^{\circ}$, and $\eta\approx-0.05$ (where both conservative and non-conservative interactions are present), as shown in Fig. 4(b). When the phase difference reaches $\Delta\varphi=\pi$, the conservative interactions dominate, resulting in non-degenerate normal modes with typical mode splitting. However, when the phase difference reaches $\Delta\varphi=1.39\pi$, approaching $\Delta\varphi=3\pi/2$, the non-conservative interactions dominate, causing two of the normal modes to become degenerate. It results in resonance and a significant amplification in the amplitude of particle motion. In our experiment, the presence of non-conservative interactions makes it difficult to observe normal mode splitting and may even lead to particle escape.
@@ 5 Position-dependence of interactions
Figure 5: (a) The normal mode splitting $|\omega_{2}-\omega_{1}|$ as a function of interparticle distance $d$ is shown for $\theta=45^{\circ}$, $\eta\approx-0.24$, and $\Delta\varphi=0$. The blue dots represent the experimentally measured normal mode splitting, while the red curve depicts the simulation results incorporating the intrinsic mechanical frequencies $\Omega_{1}$, $\Omega_{2}$, and $\Omega_{3}$ for varying distances. The curve exhibits a periodicity of $\sim\sqrt{2}\lambda/2$. The avoided crossing is unresolved when the splitting is smaller than the mechanical linewidth $\gamma$ (indicated by the dashed gray line). (b-e) The PSD of the $z$-mode signals is presented for interparticle distances $d=3.01\ \mu\rm{m}$, $3.43\ \mu\rm{m}$, $3.81\ \mu\rm{m}$, and $4.24\ \mu\rm{m}$ with $\theta=45^{\circ}$, $\eta\approx-0.24$, and $\Delta\varphi=0$. At these distances, the normal modes $\omega_{1}$ and $\omega_{2}$ exhibit variations between maximum splitting and no splitting. (f) The normal mode splitting $|\omega_{2}-\omega_{1}|$ as a function of the angle $\theta$ is shown for $d=3.01\ \mu\text{m}$, $\eta\approx-0.05$, and $\Delta\varphi=0$. The magenta dots represent the experimentally measured normal mode splitting, while the red curve illustrates the simulation results. As the angle $\theta$ varies from $35.4^{\circ}$ to $60.9^{\circ}$, $|\omega_{2}-\omega_{1}|$ oscillates, with both the period and amplitude of the oscillation gradually decreasing. The gray dashed line represents the mechanical linewidth $\gamma$.